Spike averaged 45 mph driving from Rochester to Syracuse and 49 mph driving from Syracuse to Albany. If he drove a total of 237 miles in 5 hours, then how far is it from Syracuse to Albany?
Answer by alicealc(293)

Let:
time to travel from Rochester to Syracuse = x
time to travel from Syracuse to Albany = y
the equation system will be:
x + y = 5
45x + 49y = 237
multiply the first equation by 45, we get:
45x + 45y = 225
use the elimination process:
45x + 45y = 225
45x + 49y = 237 -
---------------
-4y = -12
y = 3
then calculate the distance from Syracuse to Albany:
distance = 49y = 49*3 = 147 miles